# Odd-parity longitudinal magnetoconductivity in time-reversal symmetry broken materials

### v1

@@ I Introduction
Odd-$B$ magnetoconductivity has indeed been observed in several TRS-broken systems and studied theoretically Sahani et al. (2025); Cortijo (2016); Sharma et al. (2017); Das and Agarwal (2019); Knoll et al. (2020); Xiao et al. (2020); Zyuzin (2021); Jiang et al. (2021a); Sunko et al. (2023); Albarakati et al. (2019); Wang et al. (2020); Niu et al. (2021); Fujita et al. (2015); Jiang et al. (2021b); Takiguchi et al. (2022); Moubah et al. (2014); Hotta et al. (2025). However, its microscopic origin and crystalline symmetry restrictions remain underexplored. More importantly, a comprehensive study of the odd-$B$ longitudinal magnetoconductivity across both the low-field semiclassical and high-field quantum oscillation regimes is still lacking.
@@ II Semiclassical theory of OMC
Here, we used the shorthand $\int_{n\bm{k}}\equiv\sum_{n}\int[d\bm{k}]$, and omit explicit reference to the band index and momentum in the relevant quantities. The response $\sigma^{\rm O}_{ab}$ denotes the intrinsic Hall response induced by the BC and OMM Das and Agarwal (2021), while $\sigma^{\rm L}_{ab}$ represents the Lorentz Hall response. Both contributions are antisymmetric under the interchange of spatial indices and therefore correspond to a genuine Hall response. In contrast, the odd-parity magnetoconductivity $\sigma^{\rm OMC}_{ab}$ is symmetric under $a\leftrightarrow b$ and can contribute to both longitudinal and transverse currents. For a detailed discussion on the symmetric odd-$B$ transverse magnetoconductivity, we refer to Ref. Sunko et al. (2023).
@@
Since $\sigma^{\rm OMC}_{ab}$ is finite only in TRS–broken materials, it serves as a direct signature of intrinsic TRS breaking. In experiments, however, the longitudinal magnetoresponse always contains a $B$-independent Drude term together with conventional $B^{2}$ contributions. The odd-parity magnetoconductivity compared to the conventional even-parity component can be quantified by the following ratio
@@
III Crystalline Symmetry restriction on AHC and OMC
@@
Beyond these fundamental symmetry constraints, crystallographic point group symmetries further restrict the allowed responses. Since the anomalous Hall conductivity (AHC) can also be finite only in TRS-broken systems, it is natural to ask whether AHC and OMC can coexist under the same symmetry conditions. Now, to systematically determine the crystalline symmetry constraints for OMC, we use the following tensorial form, $\sigma^{\rm OMC}_{ab}=\tau\chi_{ab;c}^{\rm OMC}B_{c}$. Here, $\chi_{ab;c}^{\rm OMC}$ is the $\mathcal{T}$-odd axial tensor associated with OMC. In contrast, the anomalous Hall conductivity $\sigma^{\rm AH}_{ab}$ is a polar tensor. In Jahn notation, $\chi^{\rm OMC}_{ab;c}$ maps to $\texttt{ae}[\rm V^{2}]V$ and $\sigma^{\rm AH}_{ab}$ to $\texttt{a}\{\rm V^{2}\}$, with a and e denoting magnetic and axial tensors; the curly brackets indicate antisymmetry in spatial indices. For a general point group operation $\cal O$}, the OMC and the AHC response tensors transform as Newnham (2005); Gallego et al. (2019)
$\chi_{a^{\prime}b^{\prime};c^{\prime}}^{\rm OMC}$ $=$ $\eta_{\cal T}{\cal O}_{a^{\prime}a}{\cal O}_{b^{\prime}b}{\cal O}_{c^{\prime}c}\chi_{ab;c}^{\rm OMC},$ (7a)
$\sigma_{a^{\prime}b^{\prime}}^{\rm AH}$ $=$ $\eta_{\cal T}{\rm det}\{{\cal O}\}{\cal O}_{a^{\prime}a}{\cal O}_{b^{\prime}b}\sigma_{ab}^{\rm AH}~.$ (7b)
@@
The resulting symmetry restrictions for $\chi_{xx;z}^{\rm OMC}$, $\chi_{yx;z}^{\rm OMC}$, and $\sigma_{yx}^{\rm AH}$ are summarized in Table 1, for common crystalline symmetry elements. Notably, we find that $\chi_{xx;z}^{\rm OMC}$ and $\sigma_{yx}^{\rm AH}$ obey identical symmetry constraints and are allowed under the same point groups, implying that longitudinal OMC and the anomalous Hall effect always coexist. A complete classification of point groups allowing finite $\chi_{xx;z}^{\rm OMC}$ is given in Table 2 of Appendix B. In contrast, $\chi_{yx;z}^{\rm OMC}$ follows distinct symmetry rules. Specifically, in systems with an in-plane mirror or a twofold in-plane rotation symmetry, the anomalous Hall effect is forbidden, but a symmetric (in spatial indices) transverse Hall response $\propto B$ can still arise. This highlights that the symmetric $B$-linear transverse magnetoresponse also provides an independent and sensitive probe of intrinsic TRS breaking Sunko et al. (2023).
Given that the longitudinal OMC component $\chi_{xx;z}^{\rm OMC}$ remains symmetry-allowed under the same conditions as the AHC, its detection in transport, therefore, provides a direct and complementary probe of intrinsic TRS breaking. Unlike the AHC, however, the longitudinal OMC manifests in the primary transport channel, making it experimentally accessible even in systems where anomalous Hall voltage or Hall angle is small or masked by extrinsic effects Liu et al. (2025); Yue and Jin (2017); Dulal et al. (2019); Nozières, P. and Lewiner, C. (1973); Liu et al. (2024); Chowdhury et al. (2024); Manna et al. (2018).
@@ IV Odd-parity magnetoconductivity in valley-polarized gapped graphene
Figure 3: (a) $B$-linear (orange) and the $B$-quadratic (green) longitudinal magnetoconductivities versus chemical potential. (b) $B$-linear conductivity $\sigma_{yx}(B)=\sigma_{yx}^{\rm O}(B)+\sigma_{yx}^{\rm L}(B)$ (blue) and $B$-quadratic conductivity $\sigma_{yx}(B^{2})$ (red) versus chemical potential. Both panels are for valley-polarized graphene with $\Delta_{0}=0.01$ eV, and $B=1$ T. (c) Temperature dependence of OMC% calculated using the valley polarization of Eq. (11). Here, $\sigma^{\rm even}_{xx}$ includes both the Drude and $\sigma_{xx}(B^{2})$ contributions. The inset shows the magnetic order parameter $\Delta(T)$ from Eq. (11) for $\xi=+1$. We have used $\beta=0.20$, $\mu=0.02$ eV, and $T_{c}=25$ K. The OMC% follows $\Delta(T)$ and vanishes at $T_{c}$, thereby serving as an effective order parameter for TRS breaking in transport experiments.
@@ V OMC in quantizing magnetic field
Here, $\lambda=\pm 1$ represents the band index for conduction and valence bands, $n$ denotes the Landau levels indices, and $\omega_{c}=v_{F}/l_{B}$ is the cyclotron frequency with $l_{B}=\sqrt{\hbar/eB}$ being the magnetic length. The corresponding wave functions can be represented in terms of the Harmonic oscillator eigenstates, as detailed in Appendix C. Note that in the absence of a sublattice potential, the zeroth LL lies at zero energy and is valley-degenerate. A finite $\Delta$ breaks the valley degeneracy of the zeroth LL, where it has energy of $\varepsilon_{0}=-\Delta$ ($\varepsilon_{0}=\Delta$) for the $\xi=+1$ ($\xi=-1$) valley. In Fig. 4(a), we show the LLs for the $\xi=+1$ valley. The non-dispersing zeroth LL is marked by the red horizontal line.
@@
After a detailed calculation (see Appendix C for details), we obtain the collisional magnetoconductivity in the zero temperature limit to be
@@
We find that this contribution is always $B$-linear and remains finite even when $\xi\Delta\to 0$. Hence, the zeroth LL yields a universal $B$-linear magnetoconductivity in graphene, irrespective of time-reversal or inversion symmetry breaking. This behavior is consistent with Abrikosov’s quantum theory of linear magnetoresistance Abrikosov (1998, 2000), which describes systems with linear band dispersion under strong magnetic fields. In the ultra-quantum regime, where only the lowest Landau level is occupied, the magnetoresistance is therefore universally linear in $B$, irrespective of intrinsic TRS breaking, in systems with relativistic band dispersion.
@@ Appendix A Derivation of magnetoconductivity expressions
For brevity, we have used the notation $\int_{n,{\bm{k}}}\equiv\sum_{n}\int\frac{d^{d}k}{(2\pi)^{d}}$ for $d$-dimensional system. In the above equation, the first term is zero, representing the equilibrium current. The second term is the anomalous Hall effect. To explicitly derive the magnetoconductivities, we use the relations ${\cal D}=(1+\frac{e}{\hbar}{\bm{\Omega}}\cdot{\bm{B}})^{-1}\approx 1-\frac{e}{\hbar}{\bm{\Omega}}\cdot{\bm{B}}+\frac{e^{2}}{\hbar^{2}}({\bm{\Omega}}\cdot{\bm{B}})^{2}+\cdots$, $f_{n}^{0}(\tilde{\varepsilon})\approx f_{n}^{0}(\varepsilon)-{\bm{m}}\cdot{\bm{B}}(\partial_{\varepsilon}f_{n}^{0})+\frac{{\left({\bm{m}}\cdot{\bm{B}}\right)}^{2}}{2}(~\partial^{2}_{\varepsilon}f_{n}^{0})+\cdots$. Using these, the current density up to second order in magnetic field can be derived from the following expression
@@
In Eq. (20), the first term is the Drude conductivity, whereas the second term represents the Berry curvature induced anomalous Hall conductivity. In the conductivity expressions, the terms with Levi-Civita tensor and $\propto({\bm{v}}^{0}\times\bm{B})$ only contribute to the transverse conductivity, as they are antisymmetric under the exchange of $a$ and $b$. All other terms, including the term $\propto({\bm{v}}^{0}\times\bm{B})^{2}$ contribute to both longitudinal and transverse magnetoconductivities.
Appendix B Point group symmetry constraints on OMC and symmetry analysis of $\sigma_{ab}(B^{2})$
@@
$\chi_{xx;z}^{\rm OMC}$ $1,-1,2^{\prime},m^{\prime},2/m^{\prime},2^{\prime}2^{\prime}2,m^{\prime}m^{\prime}2,mm^{\prime}m^{\prime},4,4^{\prime},-4,-4^{\prime},4/m,4^{\prime}/m,42^{\prime}2^{\prime},4m^{\prime}m^{\prime},4^{\prime}m^{\prime}m,$
@@
$\sigma^{\rm AH}_{yx}$ $1,-1,2^{\prime},m^{\prime},2/m^{\prime},2^{\prime}2^{\prime}2,m^{\prime}m^{\prime}2,mm^{\prime}m^{\prime},4,-4,4/m,42^{\prime}2^{\prime},4m^{\prime}m^{\prime},-42^{\prime}m^{\prime},4/mm^{\prime}m^{\prime},3,$
@@
In Table 2, we list all magnetic point groups under which finite $\chi_{xx;z}^{\rm OMC}$, $\chi_{yx;z}^{\rm OMC}$, and $\sigma_{yx}^{\rm AH}$ are symmetry-allowed. Notably, the point groups $4^{\prime},-4^{\prime}2^{\prime}m,4^{\prime},4^{\prime}/m,$ and $4^{\prime}m^{\prime}m$ permit a finite $\chi_{xx;z}^{\rm OMC}$ by symmetry, but forbid $\sigma^{\rm AH}_{yx}$. Importantly, $\chi_{xx;z}^{\rm OMC}$ originates from Berry curvature and orbital magnetic moment. Both quantities identically vanish under the combined symmetry ${\cal C}_{4z}{\cal T}$, which is present in all of these point groups. Thus, despite being symmetry-allowed at the tensor level, a microscopic evaluation yields zero $\chi_{xx;z}^{\rm OMC}$. Thus, we conclude that longitudinal OMC and the anomalous Hall conductivity always appear together.
@@
The transverse magnetoconductivity $\propto B^{2}$ contains terms with $\tau$-scaling of $\tau^{0}$, $\tau$ and $\tau^{2}$ and $\tau^{3}$. In general, the response can be represented by
@@
Appendix C Calculation of magnetoconductivity in the presence of Landau levels

### v3

@@ I Introduction
Odd-$B$ magnetoconductivity has indeed been observed in several TRS-broken systems and studied theoretically Sahani et al. (2025); Cortijo (2016); Sharma et al. (2017); Das and Agarwal (2019a, b); Knoll et al. (2020); Xiao et al. (2020); Zyuzin (2021); Jiang et al. (2021a); Sunko et al. (2025); Albarakati et al. (2019); Wang et al. (2020); Niu et al. (2021); Fujita et al. (2015); Jiang et al. (2021b); Takiguchi et al. (2022); Moubah et al. (2014); Hotta et al. (2025). However, its microscopic origin and crystalline symmetry restrictions remain underexplored. More importantly, a comprehensive study of the odd-$B$ longitudinal magnetoconductivity across both the low-field semiclassical and high-field quantum oscillation regimes is still lacking.
@@ II Semiclassical theory of OMC
Here, we used the shorthand $\int_{n\bm{k}}\equiv\sum_{n}\int[d\bm{k}]$, and omit explicit reference to the band index and momentum in the relevant quantities. The response $\sigma^{\rm O}_{ab}$ denotes the intrinsic Hall response induced by the BC and OMM Das and Agarwal (2021), while $\sigma^{\rm L}_{ab}$ represents the Lorentz Hall response. Both contributions are antisymmetric under the interchange of spatial indices and therefore correspond to a genuine Hall response. In contrast, the odd-parity magnetoconductivity $\sigma^{\rm OMC}_{ab}$ is symmetric under $a\leftrightarrow b$ and can contribute to both longitudinal and transverse currents. For a detailed discussion on the symmetric odd-$B$ transverse magnetoconductivity, we refer to Ref. Sunko et al. (2025). The odd-parity conductivity $\sigma^{\rm OMC}_{ab}$ receives three distinct microscopic contributions. First, the orbital magnetic moment shifts the band energy, thereby modifying the group velocity and producing the ‘velocity correction’ term. Second, the OMM-modified energy shift alters the equilibrium occupation on the Fermi surface, giving an additional contribution through the distribution function. Third, the Berry curvature alters the phase-space volume in the presence of a magnetic field, resulting in the geometric ‘phase-space’ term. These three elements together constitute the full microscopic structure of the OMC response.
@@
Since $\sigma^{\rm OMC}_{ab}$ is finite only in TRS–broken materials, it serves as a direct signature of intrinsic TRS breaking. In experiments, however, the longitudinal magnetoresponse always contains a $B$-independent Drude term together with conventional $B^{2}$ contributions. The odd-parity magnetoconductivity compared to the conventional even-parity component can be quantified by the magnitude of the following ratio
@@
It is also important to understand how OMC% scales with the scattering lifetime. The odd-parity longitudinal magnetoconductivity (OMC) scales linearly with the relaxation time $\tau$, whereas the Drude contribution and the even-$B$ longitudinal conductivity scale as $\tau$ and $\tau^{3}$, respectively (see Appendix A). The relative measure $\rm OMC$$\%$ therefore scales approximately as $\tau^{-2}$. Thus, while ultra-clean samples exhibit a larger absolute OMC signal, their relative OMC% is reduced, whereas samples with moderate disorder (intermediate $\tau$) provide more favourable conditions for observing a pronounced OMC%. We note that this argument is based on the relaxation-time approximation: additional extrinsic processes, such as skew scattering, side-jump mechanisms, etc., may introduce new scattering time-dependent odd-$B$ responses Xiao et al. (2020); Sekine and MacDonald (2018). However, such processes do not alter the underlying symmetry constraints or the leading $\tau$-scaling.
III Crystalline Symmetry restriction on OMC
@@
Beyond these fundamental symmetry constraints, crystallographic point group symmetries further restrict the allowed responses. Since the anomalous Hall conductivity (AHC) can also be finite only in TRS-broken systems, it is natural to ask whether AHC and OMC can coexist under the same symmetry conditions. Now, to systematically determine the crystalline symmetry constraints for OMC, we use the following tensorial form, $\sigma^{\rm OMC}_{ab}=\tau\chi_{ab;c}^{\rm OMC}B_{c}$. Here, $\chi_{ab;c}^{\rm OMC}$ is the $\mathcal{T}$-odd axial tensor associated with OMC. The anomalous Hall conductivity $\sigma^{\rm AH}_{ab}$ is a polar tensor. In Jahn notation, $\chi^{\rm OMC}_{ab;c}$ maps to $\texttt{ae}[\rm V^{2}]V$ and $\sigma^{\rm AH}_{ab}$ to $\texttt{a}\{\rm V^{2}\}$, with a and e denoting magnetic and axial tensors; the curly brackets indicate antisymmetry in spatial indices. For a general point group operation $\cal O$}, the OMC and the AHC response tensors transform as Newnham (2005); Gallego et al. (2019)
$\chi_{a^{\prime}b^{\prime};c^{\prime}}^{\rm OMC}$ $=$ $\eta_{\cal T}{\rm det}\{{\cal O}\}{\cal O}_{a^{\prime}a}{\cal O}_{b^{\prime}b}{\cal O}_{c^{\prime}c}\chi_{ab;c}^{\rm OMC},$ (7a)
$\sigma_{a^{\prime}b^{\prime}}^{\rm AH}$ $=$ $\eta_{\cal T}{\cal O}_{a^{\prime}a}{\cal O}_{b^{\prime}b}\sigma_{ab}^{\rm AH}~.$ (7b)
@@
The resulting symmetry restrictions for $\chi_{xx;z}^{\rm OMC}$, $\chi_{yx;z}^{\rm OMC}$, and $\sigma_{yx}^{\rm AH}$ are summarized in Table 1, for common crystalline symmetry elements. Notably, we find that $\chi_{xx;z}^{\rm OMC}$ and $\sigma_{yx}^{\rm AH}$ obey identical symmetry constraints and are allowed under the same point groups, implying that longitudinal OMC and the anomalous Hall effect always coexist. A complete classification of point groups allowing finite $\chi_{xx;z}^{\rm OMC}$ and $\sigma^{\rm AH}_{yx}$ is given in Table 2 of Appendix B. We mention that although $\sigma_{xx}^{\rm OMC}$ and $\sigma^{\rm AH}_{yx}$ obey the same symmetry constraints, they appear in transport measurements in different channels. The OMC modifies the longitudinal conductivity and is odd under $B\to-B$, allowing it to be isolated by antisymmetrizing the longitudinal resistivity. In contrast, the anomalous Hall response contributes to the transverse response and is obtained from the $B=0$ offset of the Hall resistivity. These two effects, therefore, do not mix in experiments and can be independently extracted.
Because $\chi_{xx;z}^{\rm OMC}$ is symmetry-allowed under the same conditions as the AHE, detecting OMC provides a direct and complementary probe of intrinsic time-reversal-symmetry breaking. Unlike the AHC, however, the longitudinal OMC manifests in the primary transport channel, making it experimentally accessible even in systems where anomalous Hall voltage or Hall angle is small or masked by extrinsic effects Liu et al. (2025); Yue and Jin (2017); Dulal et al. (2019); Nozières, P. and Lewiner, C. (1973); Liu et al. (2024); Chowdhury et al. (2024); Manna et al. (2018).
The transverse odd-parity response $\chi_{yx;z}^{\rm OMC}$ follows distinct symmetry rules. Specifically, in systems with an in-plane mirror or a twofold in-plane rotation symmetry, the anomalous Hall effect is forbidden, but a symmetric (in spatial indices) transverse Hall response $\propto B$ can still arise. This highlights that the symmetric $B$-linear transverse magnetoresponse also provides an independent and sensitive probe of intrinsic TRS breaking Sunko et al. (2025).
@@ IV Odd-parity magnetoconductivity in valley-polarized gapped graphene
Figure 3: (a) $B$-linear (orange) and the $B$-quadratic (green) longitudinal magnetoconductivities versus chemical potential. (b) $B$-linear conductivity $\sigma_{yx}(B)=\sigma_{yx}^{\rm O}(B)+\sigma_{yx}^{\rm L}(B)$ (blue) and $B$-quadratic conductivity $\sigma_{yx}(B^{2})$ (red) versus chemical potential. Both panels are for valley-polarized graphene with $\Delta_{0}=0.01$ eV, and $B=1$ T. In both (a) and (b), the solid line represents the numerically evaluated result, whereas the open circles denote the analytical results. (c) Numerically calculated temperature dependence of OMC%, assuming valley polarization of Eq. (11). Here, $\sigma^{\rm even}_{xx}$ includes both the Drude and $\sigma_{xx}(B^{2})$ contributions. The inset shows the magnetic order parameter $\Delta(T)$ in eV of Eq. (11) for $\xi=+1$. We have used $\beta=0.20$, $\mu=15$ meV, and $T_{c}=25$ K. The OMC% follows $\Delta(T)$ and vanishes at $T_{c}$, thereby serving as an effective order parameter for TRS breaking in transport experiments.
@@ V OMC in quantizing magnetic field
Here, $\lambda=\pm 1$ represents the band index for conduction and valence bands, $n$ denotes the Landau levels indices, and $\omega_{c}=v_{F}/l_{B}$ is the cyclotron frequency with $l_{B}=\sqrt{\hbar/eB}$ being the magnetic length. The corresponding wave functions can be represented in terms of the Harmonic oscillator eigenstates, as detailed in Appendix D. Note that in the absence of a sublattice potential, the zeroth LL lies at zero energy and is valley-degenerate. A finite $\Delta$ breaks the valley degeneracy of the zeroth LL, where it has energy of $\varepsilon_{0}=-\Delta$ ($\varepsilon_{0}=\Delta$) for the $\xi=+1$ ($\xi=-1$) valley. In Fig. 4(a), we show the LLs for the $\xi=+1$ valley. The non-dispersing zeroth LL is marked by the red horizontal line.
@@
After a detailed calculation (see Appendix D for details), we obtain the collisional magnetoconductivity in the zero temperature limit to be
@@
We find that this contribution is always $B$-linear and remains finite even when $\xi\Delta\to 0$. Hence, the zeroth LL yields a universal $B$-linear magnetoconductivity in graphene, regardless of whether time-reversal or inversion symmetry is broken. This behavior is consistent with Abrikosov’s quantum theory of linear magnetoresistance Abrikosov (1998, 2000), which describes systems with linear band dispersion under strong magnetic fields. In the ultra-quantum regime, where only the lowest Landau level is occupied, the magnetoresistance is therefore universally linear in $B$, irrespective of intrinsic TRS breaking, in systems with relativistic band dispersion.
@@ Appendix A Derivation of magnetoconductivity expressions
For brevity, we have used the notation $\int_{n,{\bm{k}}}\equiv\sum_{n}\int\frac{d^{d}k}{(2\pi)^{d}}$ for $d$-dimensional system. In the above equation, the first term is zero, representing the equilibrium current. The second term is the anomalous Hall effect. To explicitly derive the magnetoconductivities, we use the relations ${\cal D}=(1+\frac{e}{\hbar}{\bm{\Omega}}\cdot{\bm{B}})^{-1}\approx 1-\frac{e}{\hbar}{\bm{\Omega}}\cdot{\bm{B}}+\frac{e^{2}}{\hbar^{2}}({\bm{\Omega}}\cdot{\bm{B}})^{2}+\cdots$, $f_{n}^{0}(\tilde{\varepsilon})\approx f_{n}^{0}(\varepsilon)-{\bm{m}}\cdot{\bm{B}}(\partial_{\varepsilon}f_{n}^{0})+\frac{{\left({\bm{m}}\cdot{\bm{B}}\right)}^{2}}{2}(~\partial^{2}_{\varepsilon}f_{n}^{0})+\cdots$. Using these, the current density up to second order in the magnetic field can be derived from the following expression
@@
In Eq. (20), the first term is the Drude conductivity, whereas the second term represents the Berry curvature-induced anomalous Hall conductivity. In the conductivity expressions, the terms with Levi-Civita tensor and $\propto({\bm{v}}^{0}\times\bm{B})$ only contribute to the transverse conductivity, as they are antisymmetric under the exchange of $a$ and $b$. All other terms, including the term $\propto({\bm{v}}^{0}\times\bm{B})^{2}$ contribute to both longitudinal and transverse magnetoconductivities.
Appendix B Point groups allowing for OMC and symmetry analysis of $\sigma_{ab}(B^{2})$
@@
$\chi_{xx;z}^{\rm OMC}$ $1,-1,2^{\prime},m^{\prime},2^{\prime}/m^{\prime},2^{\prime}2^{\prime}2,m^{\prime}m^{\prime}2,mm^{\prime}m^{\prime},4,4^{\prime},-4,-4^{\prime},4/m,4^{\prime}/m,42^{\prime}2^{\prime},4m^{\prime}m^{\prime},4^{\prime}m^{\prime}m,$
@@
$\sigma^{\rm AH}_{yx}$ $1,-1,2^{\prime},m^{\prime},2^{\prime}/m^{\prime},2^{\prime}2^{\prime}2,m^{\prime}m^{\prime}2,mm^{\prime}m^{\prime},4,-4,4/m,42^{\prime}2^{\prime},4m^{\prime}m^{\prime},-42^{\prime}m^{\prime},4/mm^{\prime}m^{\prime},3,$
@@
In Table 2, we list all magnetic point groups under which finite $\chi_{xx;z}^{\rm OMC}$, $\chi_{yx;z}^{\rm OMC}$, and $\sigma_{yx}^{\rm AH}$ are symmetry-allowed. Notably, the point groups $4^{\prime},-4^{\prime}2^{\prime}m,4^{\prime},4^{\prime}/m,$ and $4^{\prime}m^{\prime}m$ permit a finite $\chi_{xx;z}^{\rm OMC}$ by symmetry, but forbid $\sigma^{\rm AH}_{yx}$. Importantly, $\chi_{xx;z}^{\rm OMC}$ originates from Berry curvature (BC) and orbital magnetic moment (OMM). Both quantities vanish under the combined symmetry ${\cal C}_{4z}{\cal T}$, which is present in all of these point groups. Thus, despite being symmetry-allowed at the tensor level, a microscopic evaluation yields zero $\chi_{xx;z}^{\rm OMC}$. This establishes that longitudinal OMC and the anomalous Hall conductivity (AHC) always appear together.
@@
The transverse magnetoconductivity $\propto B^{2}$ contains terms with $\tau$-scaling of $\tau^{0}$, $\tau$, $\tau^{2}$ and $\tau^{3}$. In general, the response can be represented by
@@
Appendix C Temperature dependence of OMC%
The OMC% is defined as the magnitude of the ratio $\sigma_{xx}^{\rm OMC}/\sigma_{xx}^{\rm even}$. By symmetry, an odd-in-$B$ longitudinal response is permitted only when time-reversal symmetry (TRS) is broken. Consequently, its leading dependence must be proportional to the TRS-breaking parameter in the Hamiltonian. In contrast, $\sigma_{xx}^{\rm even}$ is a TRS-even quantity dominated by the Drude contribution and the conventional $B^{2}$ magnetoconductivity, both of which remain finite in TRS-preserving materials. Also, these quantities should weakly depend on the TRS-breaking parameter in the Hamiltonian. Consequently, when the temperature dependence of the TRS-breaking parameter is considered, the $\rm OMC$$\%$ inherits its temperature dependence primarily from the $\sigma^{\rm OMC}_{xx}$, which is directly proportional to the TRS-breaking parameter in leading order. Hence, OMC% closely follows the temperature evolution of the order parameter. This argument is applicable near $T\simeq T_{c}$ where the TRS-breaking parameter is small: the odd-parity response vanishes as $T$ approaches $T_{c}$, while the even-$B$ background remains minimally affected. As a result, OMC% provides a direct and experimentally accessible measure of the underlying TRS-breaking amplitude.
Having explained why OMC% should directly depend on the $\Delta(T)$ from the symmetry-based qualitative argument, we now show this explicitly using the valley polarized gapped Dirac model considered in the main text. In the expression of OMC% of Eq. (4), the $\sigma^{\rm tot}_{ab}$ is the measured total conductivity and $\sigma^{\rm even}_{ab}$ is the even-$B$ part, including the zero-field Drude contribution. This experimentally accessible ratio removes the dominant Drude and even-$B$ backgrounds, isolating the odd-parity response. To estimate OMC, we use the expression of $\sigma^{\rm even}_{xx}=\sigma^{\rm D}_{xx}+\sigma_{xx}(B^{2})$, with $\sigma^{\rm D}_{xx}=\frac{e^{2}\tau}{\hbar}\frac{\varepsilon^{2}-\Delta^{2}}{|\varepsilon|}$ being the Drude conductivity. Additionally, we have $\sigma^{\rm OMC}_{xx}=-{\rm sgn}(\varepsilon)\xi\Delta\frac{\tau e^{3}v_{F}^{2}}{4\pi\hbar}\frac{2\varepsilon^{2}-\Delta^{2}}{\varepsilon^{4}}B$ [Eq. (9a)] and $\sigma_{xx}(B^{2})=-\frac{e^{4}v_{F}^{4}}{4\pi}\frac{\tau^{3}}{\hbar^{2}}\frac{\varepsilon^{2}-\Delta^{2}}{|\varepsilon|^{3}}B^{2}$ [Eq. (10a)]. We have ignored the linear in $\tau$ contributions to $\sigma_{xx}(B^{2})$, since this originates from the band geometric quantities, which are very small compared to the conventional classical contribution $\propto\tau^{3}$. We obtain the OMC to be (valley contributions are summed up)
$\eta(\varepsilon)={\rm OMC}=e\hbar v_{F}^{2}B\frac{(2\varepsilon^{2}-\Delta^{2})}{(|\varepsilon|^{2}-C^{2})(\varepsilon^{2}-\Delta^{2})}\frac{\Delta}{|\varepsilon|},$ (25)
where $C=\tau ev_{F}^{2}B\sim\rm 4$ meV, on using $\tau=10^{-12}$ s, $v_{F}=2\times 10^{5}$ m/s, and $B=0.1$ T. For low-energy Dirac bands where $\varepsilon$ is typically around tens of meV, $C<\varepsilon$. Furthermore, $\Delta(T)$ is small around the critical temperature, assuming the chemical potential is away from the band edges, we have $\varepsilon\gg\Delta$. Under the experimentally relevant hierarchy $\Delta\ll|\varepsilon|$ and ${C}<\varepsilon$, we obtain the leading order expression of $\eta(\varepsilon)$ to be
$\eta(\varepsilon)=e\hbar v_{F}^{2}B\frac{(2\varepsilon^{2}-\Delta^{2})}{(|\varepsilon|^{2}-C^{2})(\varepsilon^{2}-\Delta^{2})}\frac{\Delta}{|\varepsilon|}\approx{2e\hbar v_{F}^{2}B}\frac{\Delta}{|\varepsilon|^{3}}+{\cal O}(\Delta^{3})~.$ (26)
In the Dirac model, the energy-resolved kernel entering the $\eta$ is $\propto\frac{\Delta}{|\varepsilon|^{3}}$ to leading order in the TRS-breaking parameter. The finite–temperature $\eta$ is obtained from the energy-resolved kernel $\eta(\varepsilon,\Delta(T))$ by the convolution
$\eta(\mu,T)=\int d\varepsilon\;[-\partial_{\varepsilon}f(\varepsilon,\mu,T)]\;\eta(\varepsilon,\Delta(T)),$ (27)
where $f^{0}$ is the Fermi–Dirac distribution function and $\Delta(T)$ is the temperature-dependent order parameter. Using the Sommerfeld regime $k_{B}T\ll|\mu|$, we have
$\eta(\mu,T)={2e\hbar v_{F}^{2}B}\int d\varepsilon\;(-\partial_{\varepsilon}f^{0})\,\frac{\Delta(T)}{\varepsilon^{3}}={2e\hbar v_{F}^{2}B}\,\Delta(T)\left[\frac{1}{\mu^{3}}+\frac{\pi^{2}}{6}(k_{B}T)^{2}\left.\frac{d^{2}}{d\varepsilon^{2}}\left(\frac{1}{\varepsilon^{3}}\right)\right|_{\varepsilon=\mu}+{\cal O}(T^{4})\right].$ (28)
This implies that the dominant contribution to the OMC% is
${\rm OMC}\%\approx 100\times\frac{2e\hbar v_{F}^{2}B}{|\mu|^{3}}\Delta(T)~.$ (29)
This clearly demonstrates that the dominant temperature dependence of the OMC% originates from $\Delta(T)$, and thermal smearing only produces small relative corrections of order $(k_{B}T/\mu)^{2}$.
Both symmetry considerations and explicit analytical evaluation show that ${\rm OMC}\%\propto\Delta(T)$ for $T\sim T_{c}$. The odd-parity signal, therefore, tracks the temperature evolution of the TRS-breaking order parameter with only weak thermal corrections.
Appendix D Calculation of magnetoconductivity in the presence of Landau levels